A 30,000 kg railroad car initially traveling to the right at 15 m/s collides inelastically with a 20,000 kg railroad car initially traveling to the left at 5 m/s. The cars stick together. What is their final speed?
1
Expert's answer
2012-11-26T09:44:55-0500
p1=30000*15
p2=20000*5
p1-p2=450000-100000=350000
p1-p2=p0
p0=m*V=>V=p0/m=350000/50000=7m/s